a car travelling at 60km/h overtakes another car travelling at 42km/h. assuming each car to be 5.0m long,find the time taken during the overtake and the total road distance used for overtake.

Grade:12

2 Answers

Vikas TU
10 years ago

a car traveling at 60 km/hr overtakes another car traveling at 42 km/hr.
Assuming each car to be 5.0 m long, find the time taken during the overtake and the total distance used for the overtake
:
Find the time (in seconds) required to pass the car
Change 5 m to km: 5/1000 = .005 km is the length of the car


:
time = .005%2F%2860-42%29 * 3600 = 10 seconds
:
Find the distance required to accomplish the pass
10%2F3600 * 60 * 1000 = 1662%2F3 meters''

 

 

 

 

Approve!

chukkaiwar rohit
10 years ago

given the 1st car travelling with 60kmph=50/3 m/s ,velocity of second car =42kmph=35/3 m/s

now length of each car =5m=s/2=>s=10m

in solving this prob we have to use concept of relative velocity i.e., rel. velocity of 1st car w.r.t. 2nd car = v1-v2=50/3-35/3=5m/s

t=s/v=10/5=2seconds

now distance used= distance travelled by 1st car during 2 seconds, s=vt=50/3x2=33.3m